OpenStudy (anonymous):

2000 people is attending to a concert if tickets cost $40. expenses are $8 per person in attendance for each concert. For each $2 of decrease in the ticket cost 200 more people will attend. calculate the price of the ticket that produces more profit

OpenStudy (amistre64):

profit = revenue - cost

OpenStudy (amistre64):

cost(x) = 8x revenue(x) = x(200/-2(x)) = x(-100x) if i did it right

OpenStudy (amistre64):

i tend to get the revenue mixed up its price * number sold

OpenStudy (amistre64):

40,2000 with a slope of -100 y = -100x +100(40)+2000 y = -100x + 6000 = number of people in attendance revenue = (-100x + 6000)x = -100x^2 + 6000x profit = revenue - cost = -100x^2 +6000x - (8(-100x + 6000)) = -100x^2 +6000x -(-800x + 48000) = -100x^2 +6000x +800x -48000 = -100x^2 +6800x -48000 thats better

OpenStudy (amistre64):

we just determine the vertex of this parabola and that is the price per ticket that brings in the most money